# Reciprocal theorem for calculating the flow rate of oscillatory channel flows

### v2

@@ IV Oscillatory flow in a straight rigid channel
$\Delta P_{a}\hat{Q}$ $=\Delta\hat{P}Q_{a}+\mathrm{i}{\text{Wo}}^{2}\int_{0}^{1}\int_{0}^{1}\hat{V}_{z}^{2}\,\mathrm{d}Y\mathrm{d}Z$ (14)
$=\Delta\hat{P}Q_{a}+\mathrm{i}\frac{1}{120}(\Delta\hat{P})^{2}{\text{Wo}}^{2}.$
Finally, substituting $\hat{Q}$ from Eq. (13) and imposing $\Delta P_{a}=\Delta\hat{P}=1$ (consistent with our pressure-drop-controlled regime), we obtain
$Q_{a}=\frac{1}{12}\left(1-\mathrm{i}\frac{1}{10}{\text{Wo}}^{2}\right).$ (15)
Oscillatory flow in a straight 2D channel has an exact solution, which allows us to ascertain the validity of our result in Eq. (15). Therefore, we review this exact solution. This flow is unidirectional such that $\boldsymbol{V}=V_{z}(Y,T)\boldsymbol{e}_{z}$ [27]. Introducing the phasors from Eq. (2) and the dimensionless variables from Eq. (8), Eqs. (3) reduce to
$\frac{\partial^{2}V_{z,a}}{\partial Y^{2}}-\frac{\partial P_{a}}{\partial Z}={\text{Wo}}^{2}\mathrm{i}V_{z,a},\qquad\frac{\partial P_{a}}{\partial Y}=0,$ (16)
subject to no slip along the channel walls, $V_{z,a}(Y=0)=V_{z,a}(Y=1)=0$. The flow is driven by a time-harmonic, oscillatory pressure distribution with the gage pressure fixed at $Z=1$. Since the flow is unidirectional, and $V_{z,a}$ does not depend on $Z$, while $P_{a}$ does not depend on $Y$, it follows that [27]:
$\frac{\partial P_{a}}{\partial Z}=\Delta P_{a}\qquad\Longleftrightarrow\qquad P_{a}(Z)=\Delta P_{a}(1-Z).$ (17)
@@
$V_{z,a}(Y)=\frac{1}{\mathrm{i}{\text{Wo}}^{2}}\left[1-\frac{\cos\left(\mathrm{i}^{3/2}(1-2Y){\text{Wo}}/2\right)}{\cos\left(\mathrm{i}^{3/2}{\text{Wo}}/2\right)}\right]{\Delta P_{a}}.$ (18)
This form of the solution and its Wo expansion is also given by Panton [29, Sec. 11.7] (see also [30, §24] for a similar but not identical discussion). From Eq. (18), the dimensionless volumetric flow rate’s amplitude is
$Q_{a}=\int_{0}^{1}V_{z,a}(Y)\,\mathrm{d}Y=\mathfrak{f}_{0}({\text{Wo}}){\Delta P_{a}},$ (19)
@@
$\mathfrak{f}_{0}({\text{Wo}})$ $\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{\mathrm{i}{\text{Wo}}^{2}}\left[1-\frac{1}{\mathrm{i}^{3/2}{{\text{Wo}}}/2}\tan\big{(}\mathrm{i}^{3/2}{\text{Wo}}/2\big{)}\right]$ (20)
@@
Fig. 2 shows the real and imaginary parts of $\mathfrak{f}_{0}({\text{Wo}})$; see also the related discussion in [6, 3].
Figure 2: The real and imaginary parts of the function $\mathfrak{f}_{0}({\text{Wo}})$ from Eq. (20), which relates the complex amplitude of the dimensionless flow to the pressure drop for oscillatory flow in a straight 2D channel. Black (dark) and red (light) dashed curves represent the small- and large-Wo asymptotic expressions given in Eq. (20).
Clearly, the result in Eq. (15) is in agreement with Eq. (19) using the ${\text{Wo}}\ll 1$ expansion from Eq. (20).
V Oscillatory flow in a nonuniform rigid channel
Now, consider a nonuniform channel with a given shape, as shown in Fig. 1(b), such that $H(Z)\neq 1$. The dimensionless solution for the steady flow in this configuration, in a pressure-drop-controlled regime, can be shown (by adapting the results from, e.g., [31]) to be
$\hat{Q}=\frac{\Delta\hat{P}}{12\int_{0}^{1}H(Z)^{-3}\,\mathrm{d}Z},\qquad\hat{V}_{z}(Y,Z)=\frac{\Delta\hat{P}}{2\int_{0}^{1}H(Z)^{-3}\,\mathrm{d}Z}\left[1-\frac{Y}{H(Z)}\right]\frac{Y}{H(Z)^{2}}.$ (21)
Employing the expansion from Eq. (12) and $\hat{V}_{z}$ from Eq. (21) into Eq. (11), we have
$\Delta P_{a}\hat{Q}$ $=\Delta\hat{P}Q_{a}+\mathrm{i}{\text{Wo}}^{2}\int_{0}^{1}\int_{0}^{H(Z)}\hat{V}_{z}^{2}\,\mathrm{d}Y\mathrm{d}Z$ (22)
$=\Delta\hat{P}Q_{a}+\mathrm{i}\frac{1}{120\int_{0}^{1}H(Z)^{-3}\,\mathrm{d}Z}(\Delta\hat{P})^{2}{\text{Wo}}^{2}\int_{0}^{1}\frac{1}{H(Z)}\,\mathrm{d}Z.$
Finally, substituting $\hat{Q}$ from Eq. (21) and imposing $\Delta P_{a}=\Delta\hat{P}=1$, we obtain
$Q_{a}=\frac{1}{12\int_{0}^{1}H(Z)^{-3}\,\mathrm{d}Z}\left[1-\mathrm{i}\frac{1}{10}{\text{Wo}}^{2}\int_{0}^{1}\frac{1}{H(Z)}\,\mathrm{d}Z\right].$ (23)
Oscillatory flow in a nonuniform 2D channel under the lubrication approximation also has an analytical solution [31], which allows us to ascertain the validity of our result in Eq. (23). The lubrication approximation leads us again to Eq. (16), now subject to $V_{z,a}(Y=0)=V_{z,a}\big{(}Y=H(Z)\big{)}=0$; see, e.g., [32] for a derivation. However, the pressure gradient cannot be constant in $Z$, thus Eq. (17) becomes
$\frac{\partial P_{a}}{\partial Z}={\Delta P_{a}\mathcal{P}^{\prime}(Z)}\qquad\Longleftrightarrow\qquad P_{a}(Z)=\Delta P_{a}\mathcal{P}(Z),$ (24)
@@ V Oscillatory flow in a nonuniform rigid channel
As in Sec. IV, it is straightforward to show that
$V_{z,a}(Y,Z)=\frac{1}{\mathrm{i}{\text{Wo}}^{2}}\left[1-\frac{\cos\left(\mathrm{i}^{3/2}(1-2Y){\text{Wo}}\,H(Z)/2\right)}{\cos\left(\mathrm{i}^{3/2}{\text{Wo}}\,{H}(Z)/2\right)}\right]{\Delta P_{a}\mathcal{P}^{\prime}(Z)}.$ (25)
@@
$Q_{a}=\int_{0}^{H(Z)}V_{z,a}(Y,Z)\,\mathrm{d}Y=\frac{H(Z)}{\mathrm{i}{\text{Wo}}^{2}}\left[1-\frac{1}{\mathrm{i}^{3/2}{\text{Wo}}\,H(Z)/2}\tan\left(\mathrm{i}^{3/2}{\text{Wo}}\,H(Z)/2\right)\right]{\Delta P_{a}\mathcal{P}^{\prime}(Z)},$ (26)
@@
$Q_{a}\int_{0}^{1}\frac{\mathrm{i}{\text{Wo}}^{2}}{H(Z)}\left[1-\frac{1}{\mathrm{i}^{3/2}{\text{Wo}}\,H(Z)/2}\tan\left(\mathrm{i}^{3/2}{\text{Wo}}\,H(Z)/2\right)\right]^{-1}\,\mathrm{d}Z=\Delta P_{a}.$ (27)
One can readily verify that the result in Eq. (23) is in agreement with the small-Wo expansion of Eq. (27).
VI Conclusion
In this communication, we revisit a basic problem in fluid mechanics using the Lorentz reciprocal theorem. Specifically, we derived closed-form expressions for flow rates in straight and nonuniform rigid 2D channels conveying oscillatory flows of Newtonian fluids. Specifically, we derived a reciprocal theorem, given by Eq. (11), which relates the flow rates and pressure drops of steady and oscillatory flows (in 2D channels) to each other. For weakly oscillatory flow, we showed how the steady (Hagen–Poiseuille) flow solution can be used as the auxiliary problem in the reciprocal theorem to obtain the oscillatory flow rate up to $\mathrm{O}({\text{Wo}}^{2})$. The flow rates obtained via the reciprocal theorem were verified by comparing to the low-Womersley-number asymptotic expressions stemming from the known analytical solutions (valid at any Womersley number). To the best of our knowledge, this admittedly simple but elegant calculation has not been done in the literature, highlighting the novelty of our approach.
@@ Appendix A Flow-rate-controlled regime
For a flow-rate-controlled regime, we impose $q_{c}=q$; consequently, $v_{c}=q/h_{0}$ and $p_{c}=\mu_{f}q\ell/h_{0}^{3}$.
@@ A.1 Straight rigid channel
$\Delta\hat{P}=12\hat{Q},\qquad\hat{V}_{z}(Y)=6\hat{Q}(1-Y)Y,$ (28)
from which the calculation from Sec. IV can be redone. Specifically, imposing $Q_{a}=\hat{Q}=1$ and substituting Eq. (28) into Eq. (11) and using Eq. (12), we obtain
$\Delta P_{a}=12+\mathrm{i}\frac{6}{5}{\text{Wo}}^{2}+\mathrm{O}({\text{Wo}}^{4}),$ (29)
which is in agreement with the small-Wo expansion of ${1}/{\mathfrak{f}_{0}({\text{Wo}})}$ from Eq. (19).
@@ A.2 Nonuniform rigid channel
$\Delta\hat{P}=\int_{0}^{1}\frac{12\hat{Q}}{H(Z)^{3}}\,\mathrm{d}Z,\qquad\hat{V}_{z}(Y,Z)=6\hat{Q}\left[1-\frac{Y}{H(Z)}\right]\frac{Y}{H(Z)^{2}},$ (30)
from which the calculation from Sec. V can be redone. Specifically, imposing $Q_{a}=\hat{Q}=1$ and substituting Eq. (30) into Eq. (11) and using Eq. (12), we obtain
$\Delta P_{a}=12\int_{0}^{1}\frac{\mathrm{d}Z}{H(Z)^{3}}+\mathrm{i}\frac{6}{5}{\text{Wo}}^{2}\int_{0}^{1}\frac{\mathrm{d}Z}{H(Z)}+\mathrm{O}({\text{Wo}}^{4}),$ (31)
which is in agreement with small-Wo expansion of the left-hand side of Eq. (27).

### v3

@@ 4 Oscillatory flow in a straight rigid channel
$\Delta P_{a}\hat{Q}$ $=\Delta\hat{P}Q_{a}+\mathrm{i}\mathrm{Wo}^{2}\int_{0}^{1}\int_{0}^{1}\hat{V}_{z}^{2}\,\mathrm{d}Y\mathrm{d}Z$ (14)
$=\Delta\hat{P}Q_{a}+\mathrm{i}\frac{1}{120}(\Delta\hat{P})^{2}\mathrm{Wo}^{2}.$
Finally, substituting $\hat{Q}$ from Eq. (13) and imposing $\Delta P_{a}=\Delta\hat{P}=1$ (consistent with the pressure-drop-controlled regime), we obtain
$Q_{a}=\frac{1}{12}\left(1-\mathrm{i}\frac{1}{10}\mathrm{Wo}^{2}\right).$ (15)
Oscillatory flow in a straight 2D channel has an exact solution, which allows us to ascertain the validity of our result in Eq. (15). Therefore, we review this exact solution. This flow is unidirectional such that $\bm{V}=V_{z}(Y,T)\bm{e}_{z}$ [27]. Introducing the phasors from Eqs. (2) and the dimensionless variables from Eq. (8), Eqs. (3) reduce to
$\frac{\partial^{2}V_{z,a}}{\partial Y^{2}}-\frac{\partial P_{a}}{\partial Z}=\mathrm{Wo}^{2}\mathrm{i}V_{z,a},\quad\frac{\partial P_{a}}{\partial Y}=0,$ (16)
subject to no slip along the channel walls, $V_{z,a}(Y=0)=V_{z,a}(Y=1)=0$. The flow is driven by a time-harmonic, oscillatory pressure at the inlet with the gage pressure fixed at $Z=1$ (the outlet). Since the flow is unidirectional, and $V_{z,a}$ does not depend on $Z$, while $P_{a}$ does not depend on $Y$, it follows that [27]:
$\frac{\partial P_{a}}{\partial Z}=\Delta P_{a}\quad\Longleftrightarrow\quad P_{a}(Z)=\Delta P_{a}(1-Z).$ (17)
@@
$V_{z,a}(Y)=\frac{1}{\mathrm{i}\mathrm{Wo}^{2}}\left[1-\frac{\cos\left(\mathrm{i}^{3/2}(1-2Y)\mathrm{Wo}/2\right)}{\cos\left(\mathrm{i}^{3/2}\mathrm{Wo}/2\right)}\right]{\Delta P_{a}}.$ (18)
This form of the solution and its $\mathrm{Wo}$ expansion is also given by Panton [29, Sec. 11.7] (see also [30, §24] for a similar but not identical discussion). From Eq. (18), the dimensionless volumetric flow rate’s amplitude is
$Q_{a}=\int_{0}^{1}V_{z,a}(Y)\,\mathrm{d}Y=\mathfrak{f}_{0}(\mathrm{Wo}){\Delta P_{a}},$ (19)
@@
$\mathfrak{f}_{0}(\mathrm{Wo})$ $\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{\mathrm{i}\mathrm{Wo}^{2}}\left[1-\frac{1}{\mathrm{i}^{3/2}{\mathrm{Wo}}/2}\tan\big(\mathrm{i}^{3/2}\mathrm{Wo}/2\big)\right]$ (20)
@@
Figure 2 shows the real and imaginary parts of $\mathfrak{f}_{0}(\mathrm{Wo})$; see also the related discussion in [6, 3].
Figure 2: The real and imaginary parts of the function $\mathfrak{f}_{0}(\mathrm{Wo})$ from Eq. (20), which relates the (complex) amplitudes of the dimensionless flow rate and pressure drop for oscillatory flow in a straight 2D channel. Dark (black) and light (red) dashed curves represent the small- and large-$\mathrm{Wo}$ asymptotic expressions given in Eq. (20).
Clearly, the result in Eq. (15) is in agreement with Eq. (19) using the small-$\mathrm{Wo}$ expansion from Eq. (20).
5 Oscillatory flow in a nonuniform rigid channel
Now, consider a nonuniform channel with a given height variation $H(Z)$ in the flow-wise direction $Z$, as shown in Fig. 1(b). The dimensionless solution for the steady flow in this configuration, in the pressure-drop-controlled regime, can be shown (by adapting the results from, e.g., [31]) to be
$\hat{Q}$ $=\frac{\Delta\hat{P}}{12\int_{0}^{1}H(Z)^{-3}\,\mathrm{d}Z},$ (21a)
$\hat{V}_{z}(Y,Z)$ $=\frac{\Delta\hat{P}}{2\int_{0}^{1}H(Z)^{-3}\,\mathrm{d}Z}\left[1-\frac{Y}{H(Z)}\right]\frac{Y}{H(Z)^{2}}.$ (21b)
Employing the expansion from Eq. (12) and substituting $\hat{V}_{z}$ from Eq. (21b) into Eq. (11), we have
$\Delta P_{a}\hat{Q}$ $=\Delta\hat{P}Q_{a}+\mathrm{i}\mathrm{Wo}^{2}\int_{0}^{1}\int_{0}^{H(Z)}\hat{V}_{z}^{2}\,\mathrm{d}Y\mathrm{d}Z$ (22)
$=\Delta\hat{P}Q_{a}+\mathrm{i}\frac{1}{120\int_{0}^{1}H(Z)^{-3}\,\mathrm{d}Z}(\Delta\hat{P})^{2}\mathrm{Wo}^{2}\int_{0}^{1}\frac{1}{H(Z)}\,\mathrm{d}Z.$
Finally, substituting $\hat{Q}$ from Eq. (21a) and imposing $\Delta P_{a}=\Delta\hat{P}=1$, we obtain
$Q_{a}=\frac{1}{12\int_{0}^{1}H(Z)^{-3}\,\mathrm{d}Z}\left[1-\mathrm{i}\frac{1}{10}\mathrm{Wo}^{2}\int_{0}^{1}\frac{1}{H(Z)}\,\mathrm{d}Z\right].$ (23)
Oscillatory flow in a nonuniform 2D channel under the lubrication approximation also has an analytical solution, though it does not appear in textbooks, which allows us to ascertain the validity of our result in Eq. (23). The lubrication approximation leads us again to Eq. (16), now subject to $V_{z,a}(Y=0)=V_{z,a}\big(Y=H(Z)\big)=0$; see, e.g., [32] for a derivation. However, the pressure gradient is no longer constant in $Z$, thus Eq. (17) becomes
$\frac{\partial P_{a}}{\partial Z}={\Delta P_{a}\mathcal{P}^{\prime}(Z)}\quad\Longleftrightarrow\quad P_{a}(Z)=\Delta P_{a}\mathcal{P}(Z),$ (24)
@@ 5 Oscillatory flow in a nonuniform rigid channel
As in Sec. 4, it is straightforward to show that
$V_{z,a}(Y,Z)=\frac{1}{\mathrm{i}\mathrm{Wo}^{2}}\left[1-\frac{\cos\left(\mathrm{i}^{3/2}(1-2Y)\mathrm{Wo}\,H(Z)/2\right)}{\cos\left(\mathrm{i}^{3/2}\mathrm{Wo}\,{H}(Z)/2\right)}\right]{\Delta P_{a}\mathcal{P}^{\prime}(Z)}.$ (25)
@@
$Q_{a}=\int_{0}^{H(Z)}V_{z,a}(Y,Z)\,\mathrm{d}Y\\ =\frac{H(Z)}{\mathrm{i}\mathrm{Wo}^{2}}\left[1-\frac{1}{\mathrm{i}^{3/2}\mathrm{Wo}\,H(Z)/2}\tan\left(\mathrm{i}^{3/2}\mathrm{Wo}\,H(Z)/2\right)\right]{\Delta P_{a}\mathcal{P}^{\prime}(Z)},$ (26)
@@
$Q_{a}\int_{0}^{1}\frac{\mathrm{i}\mathrm{Wo}^{2}}{H(Z)}\frac{\mathrm{d}Z}{\left[1-\frac{1}{\mathrm{i}^{3/2}\mathrm{Wo}\,H(Z)/2}\tan\left(\mathrm{i}^{3/2}\mathrm{Wo}\,H(Z)/2\right)\right]}=\Delta P_{a}.$ (27)
One can readily verify that the result in Eq. (23) is in agreement with the small-$\mathrm{Wo}$ expansion of Eq. (27).
6 Conclusion
In this communication, we revisited a basic problem in fluid mechanics using the Lorentz reciprocal theorem. Specifically, we derived a closed-form integral relation for calculating the flow rate in rigid 2D channels, with constant or axially varying heights, conveying oscillatory flows of incompressible Newtonian fluids. Our main result is Eq. (11), which relates the flow rates and pressure drops of steady and oscillatory flows (in 2D channels) to each other. For weakly oscillatory flow, we showed how the steady (Hagen–Poiseuille) flow solution can be used as the auxiliary problem in the reciprocal theorem to obtain the oscillatory flow rate up to $\mathrm{O}(\mathrm{Wo}^{2})$. The flow rates obtained via the reciprocal theorem were verified by comparing to the low-Womersley-number asymptotic expressions stemming from the known analytical solutions (valid at any Womersley number). To the best of our knowledge, this admittedly simple but elegant calculation has not been done in the literature, highlighting the novelty of our approach.
@@ Appendix A Flow-rate-controlled regime
For a flow-rate-controlled regime, we impose $q_{c}=q$; consequently, $v_{c}=q/h_{0}$ and $p_{c}=\mu q\ell/h_{0}^{3}$.
@@ A.1 Straight rigid channel
$\Delta\hat{P}$ $=12\hat{Q}$, (28a)
$\hat{V}_{z}(Y)$ $=6\hat{Q}(1-Y)Y,$ (28b)
from which the calculation from Sec. 4 can be redone. Specifically, imposing $Q_{a}=\hat{Q}=1$ and substituting Eqs. (28) into Eq. (11) and using Eq. (12), we obtain
$\Delta P_{a}=12+\mathrm{i}\frac{6}{5}\mathrm{Wo}^{2}+\mathrm{O}(\mathrm{Wo}^{4}),$ (29)
which is in agreement with the small-$\mathrm{Wo}$ expansion of ${1}/{\mathfrak{f}_{0}(\mathrm{Wo})}$ from Eq. (19).
@@ A.2 Nonuniform rigid channel
$\Delta\hat{P}$ $=\int_{0}^{1}\frac{12\hat{Q}}{H(Z)^{3}}\,\mathrm{d}Z,$ (30a)
$\hat{V}_{z}(Y,Z)$ $=6\hat{Q}\left[1-\frac{Y}{H(Z)}\right]\frac{Y}{H(Z)^{2}},$ (30b)
from which the calculation from Sec. 5 can be redone. Specifically, imposing $Q_{a}=\hat{Q}=1$ and substituting Eqs. (30) into Eq. (11) and using Eq. (12), we obtain
$\Delta P_{a}=12\int_{0}^{1}\frac{\mathrm{d}Z}{H(Z)^{3}}+\mathrm{i}\frac{6}{5}\mathrm{Wo}^{2}\int_{0}^{1}\frac{\mathrm{d}Z}{H(Z)}+\mathrm{O}(\mathrm{Wo}^{4}),$ (31)
which is in agreement with small-$\mathrm{Wo}$ expansion of the left-hand side of Eq. (27).
CRediT authorship contribution statement
Shrihari D. Pande: Investigation, Formal analysis, Validation, Writing - Original Draft, Visualization. Evgeniy Boyko: Conceptualization, Methodology, Investigation, Formal Analysis, Writing – original draft, Writing – review & editing, Supervision, Funding acquisition. Ivan C. Christov: Conceptualization, Methodology, Investigation, Formal Analysis, Writing – original draft, Writing – review & editing, Supervision, Funding acquisition, Project administration.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
Data availability
No data was used for the research described in the article.